OpenStudy (evansanchez):

Luis needs to run at least 150 miles in order to go with the cross country team to a summer training camp. He has forty days to pre-train for the camp. Which inequality can be used to find how many miles Luis must run each day to meet this goal? 40m ≤ 150 40m < 150 40m ≥ 150 40m > 150

OpenStudy (studygurl14):

\(\large 40\geq\Large\frac{150}{x}\rightarrow \large 40x\geq 150\)

OpenStudy (studygurl14):

x = m

OpenStudy (evansanchez):

so b

OpenStudy (studygurl14):

no C